# May 2, 2024

### v1

@@ 7 Concluding remarks
Some instances of integrable cases of the Rikitake family of dynamical systems have been considered in order to illustrate this completely generic framework and to emphasise the role of Lie bialgebras and Poisson-Lie groups in dynamical systems theory. In particular, the A system is defined onto the Lie-Poisson (1+1) Poincaré algebra, whose complete classification of Lie bialgebra structures is well-known and consists in six different non-trivial cases [28]. Among them, we have considered two representative cases whose dual Lie groups $G^{\ast}$ are the ‘book’ group and the Heisenberg group. A Poisson-Lie structure deforming the Lie-Poisson Poincaré algebra can be constructed on each of these two Lie groups, and the two associated integrable deformations of the A system have been explicitly constructed and studied. We stress that, in general, the classification of Lie bialgebra structures (and, therefore, of dual Poisson-Lie group brackets) of a given Lie algebra $\mathfrak{g}$ is by no means a trivial problem, which is only fully solved for 3D (both complex and real) Lie algebras [28].
On the other hand, we have also considered the bi-Hamiltonian Rikitake system B in order to study the possibility of obtaining bi-Hamiltonian deformations under the abovementioned framework. This implies the existence of a common cocommutator map $\delta$ for the two Lie algebra structures associated to the bi-Hamiltonian structure of the B system. In this case the answer to this question turns out to be negative, and shows that the preservation of a bi-Hamiltonian structure for Poisson-Lie deformations imposes strong constraints on the formalism. Nevertheless, the particular Rikitake case AB has been shown to admit such a common Lie bialgebra map $\delta$ for the two unerlying Lie algebras, and the bi-Hamiltonian deformation can be explicitly constructed.
@@ Acknowledgements
This work has been partially supported by Agencia Estatal de Investigación (Spain) under grant PID2019-106802GB-I00/AEI/10.13039/501100011033.

### v2

@@ 7 Concluding remarks
Some instances of integrable cases of the Rikitake family of dynamical systems have been considered in order to illustrate this completely generic framework and to emphasise the role of Lie bialgebras and Poisson-Lie groups in dynamical systems theory. In particular, the A system is defined onto the Lie-Poisson (1+1) Poincaré algebra, whose complete classification of Lie bialgebra structures is well-known and consists of six different non-trivial cases [28]. Among them, we have considered two representative cases whose dual Lie groups $G^{\ast}$ are the ‘book’ group and the Heisenberg-Weyl group. A Poisson-Lie structure deforming the Lie-Poisson Poincaré algebra can be constructed on each of these two Lie groups, and the two associated integrable deformations of the A system have been explicitly constructed and studied. We stress that, in general, the classification of Lie bialgebra structures (and, therefore, of dual Poisson-Lie group brackets) of a given Lie algebra $\mathfrak{g}$ is by no means a trivial problem, which is only fully solved for 3D (both complex and real) Lie algebras [28].
On the other hand, we have also considered the bi-Hamiltonian Rikitake system B in order to study the possibility of obtaining bi-Hamiltonian deformations under the abovementioned framework. This implies the existence of a common cocommutator map $\delta$ for the two Lie algebra structures associated to the bi-Hamiltonian structure of the B system. In this case the answer to this question turns out to be negative, and shows that the preservation of a bi-Hamiltonian structure for Poisson-Lie deformations imposes strong constraints on the formalism. Nevertheless, the particular Rikitake case AB has been shown to admit such a common Lie bialgebra map $\delta$ for the two underlying Lie algebras, and the bi-Hamiltonian deformation can be explicitly constructed.
@@ Acknowledgements
This work has been partially supported by Agencia Estatal de Investigación (Spain) under grant PID2019-106802GB-I00/AEI/10.13039/501100011033. I. Gutierrez-Sagredo thanks Universidad de La Laguna where part of the work has been done for the hospitality and support.